An elementary school has 1000 students, 200 of which are enrolled in an experimental program and 800 in a regular program. The mean EOG test score for the entire school was 79 and the mean for the experimental group was 83. What was the mean for the students in the regular program?
1
Expert's answer
2013-03-06T07:01:14-0500
The mean EOG test score for the entire school equals:
m = (m1*N1+m2*N2)/(N1+N2)
m1 - the mean for the experimental group
m2 - was the mean for the students in the regular program
N1 - number of students in an experimental program
N2 - number of students in a regular program
Therefore:
m2 = (m(N1+N2) - m1*N1)/N2 = (79*1000 - 83*200)/(1000-200) = 78
Answer:& the mean for the students in the regular program was 78.